OpenStudy (anonymous):

Write an equation of a line perpendicular to y = x + 4 in slope-intercept form that passes through the point (-2, 6).

OpenStudy (anonymous):

You could use the point slope form \[ y-y _{1}=m(x-x _{1}) \] since equation of the line is perpendicular the slope will be a negative reciprocal slope of the original line. (y = mx+b is an equation of a line where m is your slope). So if you had a line that had a slope say, 3/2, and you had to find an equation of a line perpendicular to it, the slope would be -2/3. Hope that helps.